Solve the linear equations for x,y,z
2x+y-3z=5
3x-2y-2z=3
5x-3y-2=16
1
Expert's answer
2015-09-08T07:27:44-0400
2x+y-3z=5
3x-2y-2z=3
5x-3y-2=16
| 2 1 -3|
A = | 3 -2 -2|
| 5 -3 -2|
|5 |
b= |3 |
|16 |

Ax = b -> x=A^(-1)*b



| 0.1818 -1.0 0.7273|
A^(-1) = | 0.3636 -1.0 0.4545|
| -0.0909 -1.0 0.6363|
Where from

| 0.1818 -1.0 0.7273| |5|
x = | 0.3636 -1.0 0.4545|*|3|
| -0.0909 -1.0 0.6363| |16|

Solution
x= 9.5455
y= 6.0909
z= 6.7273